OpenStudy (anonymous):

I dont understand how to do this i have tried to do it like 5 times and keep getting the wrong answer.........solve each system of equations using suvstitution method......3x-11y=9;-9+33y=18?

OpenStudy (anonymous):

Is it supposed to be $$-9x+33y = 18$$ for the second equation?

OpenStudy (anonymous):

for both of them

OpenStudy (anonymous):

no... he think you missed to write it after-9

OpenStudy (anonymous):

ok im lost the equations are 3x-11y=9 and -9x+33y=18

OpenStudy (anonymous):

then can you do what i will sayin? im helping you but you type the equation... is that alright?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first divide the second equation by 3

OpenStudy (anonymous):

there is error in your equation~~ no answer after using simultaneous equation or matrix